OpenStudy (anonymous):

2. 3x+2y=7 -3x+11=y A.(6,-3) B.(6.-7) C.(-4,19/2) D.(5.-4)

OpenStudy (mathstudent55):

Add the equations together. That will eliminate x and you can solve for y easily.

OpenStudy (anonymous):

so what am substitute the x for

OpenStudy (anonymous):

am i*

OpenStudy (mathstudent55):

3x + 2y = 7 -3x + 11= y Change bottom equation to: -3x - y = -11 Now rewrite the first equation and the new version of the second equation: 3x + 2y = 7 -3x - y = -11 Now add the equations together y = -4 There is only one solution with y = -4, so it must be the correct choice.

OpenStudy (anonymous):

ok thx

OpenStudy (mathstudent55):

wlcm